Parth (parthkohli):

\[\LARGE \text{8th GRADER'S QUESTION OF THE DAY!} \] There are 90 questions on a test. You get two marks on every correct answer. Every question you do wrong deducts 1 mark. You get 60 marks, how many did you answer correctly? Please explain along. Anyone can answer, not only 8th graders.

OpenStudy (anonymous):

I'm a ninth grader, am I allowed to answer?

Parth (parthkohli):

Yep

OpenStudy (anonymous):

OK, lets say you get x questions right and y questions wrong ... 4x - y = 90

Parth (parthkohli):

Continue please

OpenStudy (savvy):

50 right and 40 wrong....

OpenStudy (anonymous):

This is a linear equation in two variables right? I'm just guessing what the next one would be ...

Parth (parthkohli):

Savvy, you're right, but you need to explain..

OpenStudy (savvy):

@pratu043 the eqn. wud be 2x - y = 60 and x + y = 90... solve them to get x=50 and y = 40

OpenStudy (anonymous):

Yeah, savvy's right, I was silly not to think of x + y = 90

Parth (parthkohli):

Savvy. Congo, you're the winner of the first 8th grader's question of the day.. Tomorrow, I'll ask a difficult one :D

OpenStudy (savvy):

@ParthKohli we get x right and y wrong x+y=90 since total question are 90 also 2x are the marks we get for x right answers and y marks are 1*y marks are deducted from 2x for y wrong answers... so the eqn. is 2x - y = 60 solve the eqn's to get x=50 and y=40

Parth (parthkohli):

Another way to do it.. 2x - (90 - x) = 60. Solve this one and you get the answer :)

OpenStudy (savvy):

ya right.....but, no offense, you are just solving the eqn's i gave you.....